# Machine Unlearning of Pre-trained Large Language Models

### v1

@@ Abstract
This study investigates the concept of the ‘right to be forgotten’ within the context of large language models (LLMs). We explore machine unlearning as a pivotal solution, with a focus on pre-trained models–a notably under-researched area. Our research delineates a comprehensive framework for machine unlearning in pre-trained LLMs, encompassing a critical analysis of seven diverse unlearning methods. Through rigorous evaluation using curated datasets from arXiv, books, and GitHub, we establish a robust benchmark for unlearning performance, demonstrating that these methods are over $10^{5}$ times more computationally efficient than retraining. Our results show that integrating gradient ascent with gradient descent on in-distribution data improves hyperparameter robustness. We also provide detailed guidelines for efficient hyperparameter tuning in the unlearning process. Our findings advance the discourse on ethical AI practices, offering substantive insights into the mechanics of machine unlearning for pre-trained LLMs and underscoring the potential for responsible AI development. [^1] Our code is available at https://github.com/yaojin17/Unlearning_LLM
@@ 1 Introduction
Despite its potential, current research on machine unlearning in the realm of LLMs has been predominantly confined to the fine-tuned model Kumar et al. (2022); Chen and Yang (2023); Maini et al. (2024). This focus has limitations, as fine-tuning models on a retained dataset is often feasible, rendering fine-tuning phase unlearning less critical. The real challenge, and our focus in this paper, is the unlearning of pre-trained LLMs. This challenge is compounded by several factors: 1) the need to adapt existing unlearning methods from other fields to pre-trained LLMs, 2) the general lack of public availability of pre-trained data used to develop LLMs, and 3) the absence of directly comparable baselines due to the exorbitant costs of retraining pre-trained LLMs.
@@
Besides, previous studies on machine unlearning in LLMs have been limited to small-scale experiments, unlearning at most 128 samples, single corpus sources, or short context lengths Jang et al. (2023); Eldan and Russinovich (2023); Yao et al. (2023). In contrast, our work unlearns thousands of chunks, each 4096 tokens in length, from a diverse range of sources across three domains, presenting a more realistic and challenging scenario. Our main contributions and findings are:
@@ 3.1 Overview
Below, we focus on those first-order approximate unlearning methods that only exploit gradient information and are often more efficient than exact unlearning and second-order designs  [^3] We present a detailed review on exact unlearning and second-order unlearning methods in Appendix A. Their general formulation is given in Eq. (3.1), which can be extended for new unlearning methods. We will discuss how to specialize it to each method revisited below and the corresponding pros and cons.
@@ 3.2.4 Gradient Ascent + Descent or KL Divergence on Retained Set
Table 2: Overall results of unlearning an open-source LLM on a subset of pre-training data (2K GitHub code repository files)
@@ 4.2 Evaluation Metrics
To circumvent this, we introduce a surrogate evaluation approach called approximate retraining, which is inspired by membership inference attacks Shokri et al. (2017); Carlini et al. (2022) that identify performance gaps between the training and unseen data. We thus hypothesize that the retrained model will exhibit consistent performance on unseen domain-specific data, albeit inferior to its performance on trained data. Given the significant imbalance between pre-training and unlearning data volumes, we expect the retrained model’s performance on unseen data within unlearned data distributions to closely align with the original (vanilla) model’s performance. Consequently, by collecting new data from the same domain as the forget set to create an ’approximate set,’ we estimate the retrained model’s performance on the forget set by the vanilla model’s performance on this approximate set. This estimation can further guide the extent of approximate unlearning, including factors such as the learning rate or optimization steps.
@@ 4.3 Model and Datasets
We conduct experiments using the open-sourced Yi-6B [^4] https://huggingface.co/01-ai/Yi-6B LLM. To rigorously assess the effectiveness of unlearning methods, we perform tests in three distinct settings: arXiv papers, GitHub code repositories, and books. Despite their public availability, these sources may still entail copyright concerns. For instance, although arXiv provides open access to preprints, in many cases, the copyright of each individual preprint remains with its authors or the rights holders. Imagine a case where the paper’s authors would like to erase their preprints from the pre-trained LLMs (“the right to be forgotten”).
@@
Figure 1:  Figures 1a, 1b and 1c are visualization of unlearning results on GitHub code across varying learning rates while optimization steps are fixed at 4.  Figures 1d, 1e and 1f are visualization of unlearning results on GitHub code across varying optimization steps while learning rate is fixed at $2\times 10^{-5}$. In Figures 1a and 1d, values below 40 are presented on a $\log_{10}$ scale, whereas values above 40 adopt a $\log_{10^{100}}$ scale. The horizontal spring-green line in Figures 1a and 1d delineates the approximate retraining baseline as the unlearning target. For Figures 1b and 1e, the scale transitions from $\log_{10}$ for values under 20 to $\log_{10^{10}}$ for values exceeding 20.
@@ 4.4 Results
The vanilla model exhibits close performance on both the forget set and unseen approximate data, with perplexity values of 3.65 and 3.69, respectively, demonstrating the model has good generalization capabilities within the arXiv domain. Compared with the vanilla model, unlearned models exhibit a slight decline in next token prediction accuracy (e.g., from 69.02% to 68.79% after gradient ascent), signifying increased difficulty in token extraction given preceding tokens as prompts. The decrease in the AUC score of MIA on unlearned models indicates that it becomes more challenging to differentiate between the forget set and unseen data, suggesting the efficacy of unlearning methods. Notably, only the model subjected to unlearning through a combination of gradient ascent and gradient descent exhibited reduced perplexity on the retain set. This outcome can be attributed to the model learning on a general dataset sampled from the same distribution as the retain set while unlearning the forget set. Furthermore, gradient ascent emerges as the sole method to enhance average accuracy across downstream tasks post-unlearning, underscoring its superiority in maintaining the model’s overall utility.
@@ 4.5 Computational efficiency analysis
Gradient Ascend (GA) $2.95\times 10^{17}$
Fine-tune with Random Label $2.95\times 10^{17}$
Fine-tuned with adversarial samples $3.93\times 10^{17}$
@@ 4.6 Ablation studies
Taking the task of unlearning GitHub code as a case study, a series of experiments are conducted to investigate the influence of learning rate and optimization steps on the unlearning outcomes, with results shown in Figure 1. For Figures 1a, 1b and 1c, we keep optimization steps fixed at 4 and vary the learning rate between $5\times 10^{-6}$ and $1\times 10^{-4}$. In Figures 1d, 1e and 1f, the learning rate is constant at $2\times 10^{-5}$, with optimization steps ranging from 1 to 32. We evaluated the unlearned model’s perplexity on both the forget and general sets, along with average performance on downstream tasks, presenting results across seven different unlearning methods for each hyperparameter configuration.
@@ 5 Related Work
We provide an overview of current research on machine unlearning, memorization, and forgetting. A more detailed version is deferred to Appendix A.
@@ 6 Conclusion
This work primarily focuses on conducting experiments with the Yi-6B model. A significant challenge arises since most LLMs do not open-source their pre-training data, making the collection of forget sets infeasible. We encourage future research to investigate the applicability of unlearning processes to other models, including those of larger sizes such as 13B or 70B, or more complicated architecture such as the mixture of experts. Also, we mainly do experiments on three pre-training data domains. Future research should aim to explore unlearning across other domains, including Wikipedia and HackerNews. Besides, our work concentrates on unlearning copyrighted content from LLMs. Future studies could expand our methodologies to address other challenges, such as unlearning biases or harmful outputs in LLMs. Furthermore, since our methods are non-convergent or may reduce model utility until convergence, the adjustment of hyperparameters becomes crucial for ideal unlearning results. While our guidelines simplify and streamline this process, we hope that future research will develop convergent methods that are less dependent on hyperparameter adjustments.
@@ Acknowledgements
The authors would thank the Yi model and its developer 01.AI for providing a sample of pre-training data, along with permission for its use and open-sourcing.
Appendix A Related Work (Full Version)
@@
Since the seminal proposal, machine unlearning has been widely studied in ML in general Xu et al. (2024) but remains rarely explored in NLP, notably generative LLMs.  Zhang et al. (2023a) discusses the challenges and implications of unlearning and other approaches to realize RTBF in LLMs. Kumar et al. (2022) propose SISA-FC and SISA-A, two extensions of SISA for classification LMs, e.g., BERT. SISA-FC only trains fully connected task layers, and SISA-A resorts to Adapters Houlsby et al. (2019) that only update a few plug-in parameters. Chen and Yang (2023) propose an efficient unlearning method via a selective teacher-student formulation for both classification and summarization tasks. They also design a fusion mechanism to merge unlearning layers for sequential data forgetting. To unlearn knowledge in generative models, Jang et al. (2023) simply perform gradient ascent on target sequences. Wang et al. (2024) presents a selective unlearning method to minimize negative impacts on unlearned model’s capabilities and proposes evaluation metrics focusing on sensitive information. Eldan and Russinovich (2023) consider a special case of unlearning the Harry Potter books from Llama2-7b. They first use a reinforced model to identify the tokens that are most related to the unlearning target and then replace idiosyncratic terms with generic ones to generate alternative labels for fine-tuning the model. Yao et al. (2023) applies machine unlearning for harmful responses removing and hallucinations eliminating. Maini et al. (2024) presents a benchmark for unlearning fictitious authors on fine-tuned models.

### v3

@@ Abstract
This study investigates the concept of the ‘right to be forgotten’ within the context of large language models (LLMs). We explore machine unlearning as a pivotal solution, with a focus on pre-trained models–a notably under-researched area. Our research delineates a comprehensive framework for machine unlearning in pre-trained LLMs, encompassing a critical analysis of seven diverse unlearning methods. Through rigorous evaluation using curated datasets from arXiv, books, and GitHub, we establish a robust benchmark for unlearning performance, demonstrating that these methods are over $10^{5}$ times more computationally efficient than retraining. Our results show that integrating gradient ascent with gradient descent on in-distribution data improves hyperparameter robustness. We also provide detailed guidelines for efficient hyperparameter tuning in the unlearning process. Our findings advance the discourse on ethical AI practices, offering substantive insights into the mechanics of machine unlearning for pre-trained LLMs and underscoring the potential for responsible AI development. [^1] Our dataset and code is available at https://github.com/yaojin17/Unlearning_LLM
@@ 1 Introduction
Figure 1: Overview of unlearning pre-trained LLMs to address user removal requests.
@@
Despite its potential, current research on machine unlearning in the realm of LLMs has been predominantly confined to the fine-tuned model Kumar et al. (2022); Chen and Yang (2023); Maini et al. (2024). While fine-tuning datasets typically contain token numbers on the magnitude of $10^{6}$ or less, pre-training corpora are on the magnitude of $10^{12}$. Consequently, retraining on the retained set of a fine-tuning dataset based on a pre-trained model is often feasible, whereas retraining the pre-trained model from scratch is impractical. Therefore, this focus on unlearning fine-tuned models has limitations and is less critical. The real challenge, and our focus in this paper, is the unlearning of pre-trained LLMs. This challenge is compounded by several factors: 1) the need to adapt existing unlearning methods from other fields to pre-trained LLMs, 2) the general lack of public availability of pre-trained data used to develop LLMs, and 3) the absence of directly comparable baselines due to the exorbitant costs of retraining pre-trained LLMs.
@@
Besides, previous studies on machine unlearning in LLMs have been limited to small-scale experiments, unlearning at most 128 samples, single corpus sources, or short context lengths Jang et al. (2023); Eldan and Russinovich (2023); Yao et al. (2023). In contrast, our work unlearns thousands of chunks, each 4096 tokens in length, from a diverse range of sources across three domains, presenting a more realistic and challenging scenario. Fig. 1 illustrates a process overview for unlearning pre-trained LLMs to address user removal requests. Our main contributions and findings are summarized below.
@@
  • We open-source our dataset, including samples from a real pre-training dataset, which is typically a private and closed-source asset.
@@ 3.1 Overview
Below, we focus on those first-order approximate unlearning methods that only exploit gradient information and are often more efficient than exact unlearning and second-order designs  [^3] We present a detailed review on exact unlearning and second-order unlearning methods in Appendix B. Their general formulation is given in Eq. (3.1), which can be extended for new unlearning methods. We will discuss how to specialize it to each method revisited below and the corresponding pros and cons.
@@ 3.2.4 Gradient Ascent + Descent or KL Divergence on Retained Set
Table 2: Overall results of unlearning an open-source LLM on a subset of pre-training data (2K GitHub code repository files). Results under multiple training random seeds are displayed in Table 6.
@@ 4.2 Evaluation Metrics
To circumvent this, we introduce a surrogate evaluation approach called approximate retraining, which is inspired by membership inference attacks Shokri et al. (2017); Carlini et al. (2022) that identify performance gaps between the training and unseen data. We thus hypothesize that the retrained model will exhibit consistent performance on unseen domain-specific data, albeit inferior to its performance on trained data. Given the significant imbalance between pre-training and unlearning data volumes, we expect the retrained model’s performance on unseen data within unlearned data distributions to closely align with the original (vanilla) model’s performance. Consequently, by collecting new data from the same domain as the forget set to create an "approximate set," we estimate the retrained model’s performance on the forget set by the vanilla model’s performance on this approximate set. This estimation can further guide the extent of approximate unlearning, including factors such as the learning rate or optimization steps. This approximate set can be easily obtained from the test set of the LLM for its developer or from the open-source community [^4] https://huggingface.co/RealTimeData for all researchers across several domains.
@@ 4.3 Model and Datasets
We conduct experiments using the open-sourced Yi-6B Young et al. (2024) LLM. To rigorously assess the effectiveness of unlearning methods, we perform tests in three distinct settings: arXiv papers, GitHub code repositories, and books. Despite their public availability, these sources may still entail copyright concerns. For instance, although arXiv provides open access to preprints, in many cases, the copyright of each individual preprint remains with its authors or the rights holders. Imagine a case where the paper’s authors would like to erase their preprints from the pre-trained LLMs (“the right to be forgotten”).
@@
Figure 2:  Figures 2a, 2b and 2c are visualization of unlearning results on GitHub code across varying learning rates while optimization steps are fixed at 4.  Figures 2d, 2e and 2f are visualization of unlearning results on GitHub code across varying optimization steps while learning rate is fixed at $2\times 10^{-5}$. In Figures 2a and 2d, values below 40 are presented on a $\log_{10}$ scale, whereas values above 40 adopt a $\log_{10^{100}}$ scale. The horizontal spring-green line in Figures 2a and 2d delineates the approximate retraining baseline as the unlearning target. For Figures 2b and 2e, the scale transitions from $\log_{10}$ for values under 20 to $\log_{10^{10}}$ for values exceeding 20.
@@ 4.4 Results
The vanilla model exhibits close performance on both the forget set and unseen approximate data, with perplexity values of 3.65 and 3.69, respectively, demonstrating the model has good generalization capabilities within the arXiv domain. Compared with the vanilla model, unlearned models exhibit a slight decline in next token prediction accuracy (e.g., from 69.02% to 68.79% after gradient ascent), signifying increased difficulty in token extraction given preceding tokens as prompts. Notably, only the model subjected to unlearning through a combination of gradient ascent and gradient descent exhibited reduced perplexity on the retain set. This outcome can be attributed to the model learning on a general dataset sampled from the same distribution as the retain set while unlearning the forget set. Furthermore, gradient ascent emerges as the sole method to enhance average accuracy across downstream tasks post-unlearning, underscoring its superiority in maintaining the model’s overall utility.
Both the MIA AUC scores before and after unlearning are close to 0.5. This result aligns with a recent study Duan et al. (2024), which indicates that current MIA methods struggle to differentiate between member and non-member data in LLMs. Despite this general difficulty, our results demonstrate a decrease in MIA AUC scores after unlearning. This indicates that it becomes more challenging to differentiate between the forget set and unseen data, suggesting that our unlearning methods are effective in reducing privacy leakage related to membership inference.
@@ 4.5 Computational efficiency analysis
Gradient Ascent (GA) $2.95\times 10^{17}$
Fine-tuning with Random Label $2.95\times 10^{17}$
Unlearning with adversarial samples $3.93\times 10^{17}$
@@ 4.6 Ablation studies
Taking the task of unlearning GitHub code as a case study, a series of experiments are conducted to investigate the influence of learning rate and optimization steps on the unlearning outcomes, with results shown in Figure 2. For Figures 2a, 2b and 2c, we keep optimization steps fixed at 4 and vary the learning rate between $5\times 10^{-6}$ and $1\times 10^{-4}$. In Figures 2d, 2e and 2f, the learning rate is constant at $2\times 10^{-5}$, with optimization steps ranging from 1 to 32. We evaluated the unlearned model’s perplexity on both the forget and general sets, along with average performance on downstream tasks, presenting results across seven different unlearning methods for each hyperparameter configuration. The discussion of the effect of batch size and forget set size on the unlearning results is deferred to Appendix A.
@@ 5 Related Work
We provide an overview of current research on machine unlearning, memorization, and forgetting. A more detailed version is deferred to Appendix B.
@@ 6 Conclusion
This work primarily focuses on conducting experiments with the Yi-6B model. A significant challenge arises since most LLMs do not open-source their pre-training data, making the collection of forget sets infeasible. We encourage future research to investigate the applicability of unlearning processes to other models, including those of larger sizes such as 13B or 70B, or more complicated architecture such as the mixture of experts. Additionally, our experiments are mainly conducted on three specific pre-training data domains. Future research should aim to explore unlearning across other domains, including Wikipedia and News.
Moreover, our work concentrates on unlearning copyrighted content from LLMs. Future studies could expand our methodologies to address other challenges, such as unlearning biases or harmful outputs in LLMs. Given that our methods are non-convergent and may reduce model utility until convergence, the adjustment of hyperparameters becomes crucial for ideal unlearning results. While our guidelines simplify and streamline this process, we hope that future research will develop convergent methods that are less dependent on hyperparameter adjustments.
This study emphasizes the practical aspects of approximate unlearning in LLMs. However, since distinguishing between member and non-member data in LLMs remains challenging, the evaluation of approximate unlearning can be complex. A more principled theoretical investigation of unlearning in LLMs is necessary for future work. Additionally, more powerful MIA methods and alternative methods beyond MIA are needed for evaluating unlearning methods in the context of LLMs. While our focus is on unlearning data within a single domain while maintaining the unlearned model’s general capability, exploring the unlearning of task-specific data and assessing the impact on the unlearned model’s performance in that task could be an interesting and important direction for future research.
@@ Acknowledgements
The authors thank the Yi model and its developer for providing a sample of pre-training data, along with permission for its use and open-sourcing. Wang was partially supported by NSF OAC-2319988.
Appendix A Additional Ablation Studies
Figures 3a, 3b and 3c display the results of unlearning 2720 sequences of GitHub code, with the batch size varied from 85 to 2720 and other hyperparameters fixed. The results demonstrate that a small batch size leads to more optimization steps, which can cause detrimental damage to the model’s general utility. However, a higher batch size also reduces the unlearning efficacy. As mentioned in Section 4.6, a batch size with corresponding optimization steps of four would be the best practice to balance utility and unlearning effect.
Figures 3d, 3e and 3f show the results of unlearning GitHub code with varying forget set sizes (512 to 8192 sequences) and fixed hyperparameters. The results show that excessively large unlearning data sizes lead to an exponential increase in perplexity on both the forget and general sets and also cause a significant decrease in model general utility. Integrating gradient ascent with gradient descent on in-distribution data proves most stable, aligning with insights in Section 4.6. For unlearning large datasets, we recommend using a relatively large batch size for stability.
(a) Forget Set PPL
(b) Sampled General Set PPL
(c) Avg. Downstream Tasks performance
(d) Forget Set PPL
(e) Sampled General Set PPL
(f) Avg. Downstream Tasks performance
Figure 3:  Figures 3a, 3b and 3c are visualization of unlearning results on GitHub code across varying batch sizes while other hyperparameters are fixed.  Figures 3d, 3e and 3f are visualization of unlearning results on GitHub code across varying forget set size while other hyperparameters are fixed. In Figures 3a and 3d, values below 40 are presented on a $\log_{10}$ scale. Values in Figure 3a above 40 adopt a $\log_{10^{100}}$ scale, whereas values in Figure 3d above 40 adopt a $\log_{10^{10}}$ scale. The horizontal spring-green line in Figures 3a and 3d delineates the approximate retraining baseline as the unlearning target. For Figures 3b and 3e, the scale transitions from $\log_{10}$ for values under 20 to $\log_{10^{10}}$ for values exceeding 20.
Models Forget Set Retain Set Downstream Task Accuracy $\uparrow$
[…]
Approximate Retrain 72.91 3.42 - - - - - - - -
Gradient Ascent 78.18 $\pm$ 0.03 3.56 $\pm$ 0.03 74.23 $\pm$ 0.05 52.6 $\pm$ 0.01 9.31 $\pm$ 0 63.47 $\pm$ 0.05 68.23 $\pm$ 0.14 14.83 $\pm$ 0.93 34.77 $\pm$ 0.39 45.33 $\pm$ 0.16
Fine-tuning with Random Labels 77.98 $\pm$ 0.02 3.12 $\pm$ 0.01 80.52 $\pm$ 0.03 52.51 $\pm$ 0.01 9.47 $\pm$ 0 62.43 $\pm$ 0.09 67.13 $\pm$ 0.1 11.39 $\pm$ 0.35 29.74 $\pm$ 0.94 42.67 $\pm$ 0.3
Unlearning with Adversarial Samples 75.17 $\pm$ 0.08 3.39 $\pm$ 0.01 79.54 $\pm$ 0.13 52.54 $\pm$ 0.01 9.41 $\pm$ 0 62.34 $\pm$ 0.05 67.41 $\pm$ 0.13 10.78 $\pm$ 0.93 31.64 $\pm$ 0.27 43.04 $\pm$ 0.3
Gradient Ascent + Descent on retain set
- Descent on in-distribution data 76.91 $\pm$ 0.03 3.44 $\pm$ 0.01 76.76 $\pm$ 0.01 52.48 $\pm$ 0.01 9.38 $\pm$ 0 62.34 $\pm$ 0.03 66.76 $\pm$ 0.02 2.85 $\pm$ 0.7 30.61 $\pm$ 0.7 40.64 $\pm$ 0.01
- Descent on general data 78.79 $\pm$ 0 3.57 $\pm$ 0.01 75.61 $\pm$ 0.01 53.03 $\pm$ 0.01 9 $\pm$ 0 63.19 $\pm$ 0.05 67.56 $\pm$ 0.06 15.04 $\pm$ 0.7 33.26 $\pm$ 0.58 44.76 $\pm$ 0.04
Gradient Ascent + KL divergence
- KL on in-distribution data 78.78 $\pm$ 0 3.51 $\pm$ 0.01 76.19 $\pm$ 0 52.61 $\pm$ 0.01 9.31 $\pm$ 0 63.44 $\pm$ 0.04 68.23 $\pm$ 0.02 14.63 $\pm$ 0 34.85 $\pm$ 0.25 45.29 $\pm$ 0.06
- KL on general data 78.68 $\pm$ 0 3.58 $\pm$ 0 75.42 $\pm$ 0 52.6 $\pm$ 0 9.31 $\pm$ 0 63.39 $\pm$ 0.06 68.01 $\pm$ 0.06 14.43 $\pm$ 0.7 35 $\pm$ 0.31 45.21 $\pm$ 0.18
Table 6: Overall results of unlearning an open-source LLM on a subset of pre-training data (2K GitHub code repository files) under random seed 42, 420, and 4200. The values are displayed in the mean $\pm$ standard deviation format.
Appendix B Related Work (Full Version)
@@
Since the seminal proposal, machine unlearning has been widely studied in ML in general Xu et al. (2024); Gong et al. (2024) but remains rarely explored in NLP, notably generative LLMs.  Zhang et al. (2023a) discusses the challenges and implications of unlearning and other approaches to realize RTBF in LLMs. Kumar et al. (2022) propose SISA-FC and SISA-A, two extensions of SISA for classification LMs, e.g., BERT. SISA-FC only trains fully connected task layers, and SISA-A resorts to Adapters Houlsby et al. (2019) that only update a few plug-in parameters. Chen and Yang (2023) propose an efficient unlearning method via a selective teacher-student formulation for both classification and summarization tasks. They also design a fusion mechanism to merge unlearning layers for sequential data forgetting. To unlearn knowledge in generative models, Jang et al. (2023) simply perform gradient ascent on target sequences. Wang et al. (2024) presents a selective unlearning method to minimize negative impacts on unlearned model’s capabilities and proposes evaluation metrics focusing on sensitive information. Eldan and Russinovich (2023) consider a special case of unlearning the Harry Potter books from Llama2-7b. They first use a reinforced model to identify the tokens that are most related to the unlearning target and then replace idiosyncratic terms with generic ones to generate alternative labels for fine-tuning the model. Yao et al. (2023) applies machine unlearning for harmful responses removing and hallucinations eliminating. Maini et al. (2024) presents a benchmark for unlearning fictitious authors on fine-tuned models.